OpenStudy (anonymous):

Medal + Fan + Testimonial! When Elyse moved last weekend, the circular glass top for an end table broke. To replace it, she bought a circular glass top with a diameter of 22 inches. What is the approximate area of the circular glass top Elyse bought?(Use 3.14 for Pi) 69.08 in2^ 1,519.76 in2^ 138.16 in2^ 379.94 in2^

OpenStudy (samsungfanboy):

@AriPotta @satellite73 @amistre64 @johnweldon1993

OpenStudy (anonymous):

It is >.<

OpenStudy (anonymous):

But nvm now thx tho :)

OpenStudy (samsungfanboy):

@mathmate @campbell_st

OpenStudy (samsungfanboy):

oh.. okay :)

OpenStudy (mathmate):

Use area = \(\pi r^2\) with r=22/2=11